OpenStudy (migitmack):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)? I did not do this in a LONG TIME

myininaya (myininaya):

So do you know how to find the equation of the line that passes through the points given?

myininaya (myininaya):

Do you know how to find the slope this line that passes through the points given?

myininaya (myininaya):

Do you know the slope formula?

myininaya (myininaya):

\[slope=m=\frac{y_2-y_1}{x_2-x_1}\] Let me know what you get for the slope when you get it Also you will need to know what form a line has One form any line can be written in is y=mx+b Where m is the slope and b is the y-intercept So once you get a known value for m You will use that value for m and use either (x=-5,y=3) or (x=1,y=-3) And plug into y=mx+b to solve for b --- Ok but one thing at a time What did you get for m?

OpenStudy (migitmack):

How di u get the slope

myininaya (myininaya):

Plug your points into the formula I gave

myininaya (myininaya):

\[(x_1=-5,y_1=3) \text{ and } (x_2=1,y_2=-3)\]

OpenStudy (anonymous):

I'm new to this, and I know Im not supposed to ask but how are u a mod cuz i wanna b 1 please

myininaya (myininaya):

Plug into \[m=\frac{y_2-y_1}{x_2-x_1}\]

myininaya (myininaya):

Hey @kkambroise12345 please talk about this in chat, message me, or post in feedback.

OpenStudy (migitmack):

ok lemme see i got -4?

OpenStudy (migitmack):

-4 is the slope?

OpenStudy (migitmack):

cuz 3-3=0 and 1-5=-4

myininaya (myininaya):

\[m=\frac{-3-3}{1-(-5)}\] I'm not seeing how to get -4 from this Try again

OpenStudy (migitmack):

-6,6

OpenStudy (migitmack):

i got -6,6 is that correct?

myininaya (myininaya):

-6/6 =?

OpenStudy (migitmack):

oh i have 2 divide too 1 seec

OpenStudy (migitmack):

-1

myininaya (myininaya):

yes m=-1 very good So we have so far y=-1x+b I just replace m with -1 since m=-1 Now we need to find b And we know a point on the line (use either point) (-5,3) x=-5,y=3 Plug this into y=-1x+b and solve for b

OpenStudy (migitmack):

i can use either -5 or 3?

myininaya (myininaya):

use -5 for x and 3 for y because x=-5 and y=3 Replace x with -5 and replace y with 3 in the eqaution y=-1x+b And then solve for b

OpenStudy (migitmack):

ok

OpenStudy (migitmack):

3=-1(-5)+b

myininaya (myininaya):

Right now we have 3=5+b solve for b :)

OpenStudy (migitmack):

3=5+b is what I have now but What now? Do i multiply 3 to 5 and 3 to b?

OpenStudy (migitmack):

WAIT NOOO

OpenStudy (migitmack):

I subtract RIGHT?!!

myininaya (myininaya):

Yes subtract what on both sides?

OpenStudy (migitmack):

OK I GOT -2

OpenStudy (migitmack):

I subtract 3 from both sides

myininaya (myininaya):

b=-2 is correct! :)

OpenStudy (migitmack):

YES WOOT FINALLY!!!

myininaya (myininaya):

so we have the equation of the line is y=-1x-2

myininaya (myininaya):

Now the last thing to do is...

OpenStudy (migitmack):

AH took 4 ever

OpenStudy (migitmack):

MORE !!

OpenStudy (migitmack):

>.<

myininaya (myininaya):

Since we want to find the x-intercept set y=0 and solve for x

OpenStudy (migitmack):

Huh

myininaya (myininaya):

Since we want to find the x-intercept, replace y with 0 and solve for x Remember we are looking at y=-1x-2

OpenStudy (migitmack):

ohh ok so uhm let me see

OpenStudy (migitmack):

2?

OpenStudy (migitmack):

-2

myininaya (myininaya):

Ok so did you replace y with 0 like this 0=-1x-2 And you want to solve this for x So how do you isolate x ?

OpenStudy (migitmack):

Uhm I think i have to subtraact

myininaya (myininaya):

Ok since you have subtraction of 2 You need to add 2 If you add 2 on one side, add 2 on the other , right?

OpenStudy (migitmack):

ok

OpenStudy (migitmack):

x=2

myininaya (myininaya):

Nope nope ok so you need to add 2 on both sides

myininaya (myininaya):

oh wait

myininaya (myininaya):

almost actually

myininaya (myininaya):

You are missing something on your answer

OpenStudy (migitmack):

I am

myininaya (myininaya):

So when you added 2 what is the exact equation you get ?

myininaya (myininaya):

After performing that one step

OpenStudy (migitmack):

x=-2

myininaya (myininaya):

Right :)

OpenStudy (migitmack):

Then what

myininaya (myininaya):

Thats it

OpenStudy (migitmack):

Wait but it does not seem right my answer is -2,-2? Or jusr -2

OpenStudy (anonymous):

You found x = -2 by setting y = 0. So the answer is (-2,0)